OpenStudy (anonymous):

help will give medal Earl needs to fill juice cups for his little brother's birthday party. There are 25 juice cups and each juice cup holds 12 fluid ounces. The juice comes in 1-gallon bottles. How many 1-gallon bottles of juice will Earl need to purchase? (1 quart = 32 fluid ounces, 1 gallon = 4 quarts) (2 points) 1 bottle 2 bottles 3 bottles

OpenStudy (breanna1999):

3 bottles

OpenStudy (anonymous):

or 4 bottles

OpenStudy (campbell_st):

start with 25 x 12 = how many fl oz you need then 1 gallon contains 32 x 4 = the divide the fl oz needed by the fl oz in 1 gallon

OpenStudy (anonymous):

First let's find out how many ounces we need. \[25 Cups * 12 Ounces per \Cup = 300 Ounces\] Next turn Ounces into gallons \[\frac{ 300 Ounces }{ 1 } * \frac{ 1 Quart }{ 32 Ounces } * \frac{ 1 Gallon }{ 4 Quarts } = 2.34375\] So in total we need 3 bottles.

OpenStudy (rushwr):

if 1quart = 32 ounces 4quarts = 32 * 4 ounces = 1 gallon right? So in 1 gallon = 128 ounces.

OpenStudy (rushwr):

i cup needs 12 ounces and u have 25 cups that makes the total fluid 25* 12 = 300 ounces If one bottle contains 1 gallon = 128ounces the no. of bottles needed = 300/128 \[\frac{ 300 }{ 128 } = 2.3 \] u need more than 2 bottles. So u would have to go with 3 bottles.

OpenStudy (rushwr):

Hope u got it @shadow2239

OpenStudy (anonymous):

ok I understand thank you could you maybe check my answer for one more question plz

OpenStudy (rushwr):

yeah

OpenStudy (anonymous):

Kenya exchanges $200 for euros (€). Suppose the conversion rate is €1 = $1.321. How many euros should Kenya receive? (1 point) €264.20 €151.40 €132.10 €66.05 my answer is A

OpenStudy (rushwr):

nop

OpenStudy (anonymous):

where did I go wrong ???

OpenStudy (rushwr):

if 1euro = $1.321 then $200 will be equal to \[\frac{ 200 }{ 1.321} = $151.4\]

OpenStudy (rushwr):

sorry 151.4 should ne in euros not in dollars

OpenStudy (anonymous):

was I suppose to divide it ?

OpenStudy (rushwr):

yeah .

OpenStudy (anonymous):

I see where I went wrong thank you so much I got an A on my test so im very happy lol now I have two more tests to go lol

OpenStudy (rushwr):

see if 1euro = $1.321 then \[$1 = \frac{ \epsilon 1 }{ 1.321}\] right? So when u need to find the value for $200 then u have to multiply it by 200 \[\frac{ 1 }{ 1.321 } *200 = \frac{ 200 }{ 1.321 }\]

OpenStudy (rushwr):

Anytym bubs :)